Fermi National Laboratory


Accelerator Update for August 31, 2005 - September 2, 2005

Wednesday, August 31
The day shift began with Operations monitoring store 4359 and the Recycler stash circulating. Accesses to MiniBooNe, NuMI, PBar, and parts of switchyard were underway. Booster was preparing for RF work and accesses. MI-12B was accessed for floor drain work.

A small leak was found on a magnet in the AP2 line and repaired.

Main Injector power supply work was performed and the replacement switchgear was installed with a crane at MI-60, but electrical connections will be made at a later date.

At 0944 a failed dehumidifier pump was found at MI-12B and needed to be replaced.

Linac worked on switching to the H- Source.

At 1020, the Pbar interlocks were dropped. Operations made plans to secure the enclosure again while people were at lunch.

Shielding work at MC6 was being done for future operations in the Meson area and was not expected to be done until 1500 today.

By 1240 the Linac, Booster, and Main Injector were back on. Startup was in progress for these machines.

Beam was established to the Main Injector dump. NuMI finished their access and had beam at 1405.

At 1430, a fault in Feeder 47 disrupted operations. It caused a quench in the Tevatron which resulted in the loss of the store. Other accelerator systems were glitched too, including the Main Injector ramp, Proton Source and Tevatron RF.

Given the opportunity with the loss of the store, CDF and D0 began controlled accesses at 1540.

By 1600 beam was restored to NuMI and within 40 minutes beam had returned to MiniBooNe.

PBar target scans began at 1730 and the Tevatron was ramping by 1820.

PBar studies revealed that the beam was not making it to the Accumulator very well and stacking problems continued through the remainder of the shift.

Collider Detector at Fermilab (CDF) and D-Zero (D0) remained in access until the end of the evening shift.

Due to the shielding work, MIPP (Meson Center) came up with reduced intensity beam at 1830.

Thursday, September 1
The midnight shift began with a nice Recycler stash, with Pbar attempting to stack, and with MiniBooNE, NuMI, and MCenter taking beam.

At 0040, Tevatron operation began and the Tevatron was in a dry squeeze by 0105.

Tevatron studies and a 36x0 store were used to checkout.

At 0420, shot setup began.

After a few problems, store 4363 was in the Tevatron and colliding by 0634. The average initial luminosity was 91.64e30.

At 0655, beam to MIPP was ended so that shielding work in MC6 could proceed during the day shift.

Since MIPP was down, M02-M05 were made ready for access by 0900.

At 0905 KRF3 tripped and would not return to full power. Experts were contacted to investigate.

After many hours of studies, the PBar rings were prepared for access at 1110. The alignment crew was asked to check the quads that were moved during yesterday's access.

A leak in the Main Injector continued, so Operations prepared the Main Injector for access at 1140.

By 1215 the shielding work in Meson Center was complete they were authorized to return operation to full intensity.

At 1315 a leak was found in the tunnel at MI-60 where some damper relief valves were poorly seated. They were reseated and the leak investigation continued.

A second leak was found in the Main Injector. This leak was on NuMI magnet, V108-5. The leak was not that bad and the time to fix it would have been longer than the available access time. It was decided to wait to repair this leak.

At 1440, the power supply for the solenoid #2 for KRF2 failed and took about a half hour to replace.

Main Injector keys were returned and turn on began at around 1500.

Beam returned to MiniBooNe and NuMI within a half hour. The investigation into PBar stacking problems also resumed at this time.

Beam was restored to switchyard at 1548 and beam returned to MIPP by 1600. Another access into the PBar rings was requested and it was configured for access at this time.

The PBar access was over within an hour and stacking investigation continued. At 1712 D:IB tripped and was emitting a bad smell. Experts were contacted to repair it.

PBar studies were completed, D:IB was repaired, and PBar returned to stacking at 2220.

NuMI tripped on a radiation detector. Intensity was lowered, things were tuned, and NuMi was back up to intensity in 15 minutes.

Friday, September 2
The midnight shift began with Operations monitoring store 4363, with Pbar stacking, and with MiniBooNE and NuMI taking beam. There isn't a Recycler stash.

Had several trips while trying to increase intensity to MiniBooNe, things were turned down and tuned.

At 0340, shots to the Recycler begin, meaning also beam was disabled to SY120. The shots were delayed due to a Recycler correction element problem.

By 0520, Recycler stash was done and beam returned to SY120.

The Plans for Friday and the Weekend (through Labor Day)
The plans for today are to begin the end-of-store studies at 1000 and they will last approximately 2 hours. After the store is terminated, CDF and D0 can access for 2 hours. A 3-4 hour Tevatron study period will follow the access termination. A new store will be put in the Tevatron. The rest of today and the weekend will consist of stacking, stashing, and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/2/2005   email Fermilab